OpenStudy (anonymous):

Determine convergence or divergence. (-1)^n/(n^3) I know that I want to use the alternating series test. I used (bsubn) as 1/n^3 and (bsubn+1) as 1/(n+1)^3. I know that (bsubn) as n ----> infinity equals 0, but is (bsubn+1) less than or equal to (bsubn)?

OpenStudy (anonymous):

Or do i need to use a comparison test?

OpenStudy (anonymous):

I know that those two conditions have to be met for it to be convergent...

OpenStudy (anonymous):

is it \[\sum_{n=1}^{\infty}\frac{(-1)^n}{n^3}\]

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

it not only converges, but it converges absolutely

OpenStudy (anonymous):

all you need for the alternating series to converge is that the terms go to zero, which thsee certainly do

OpenStudy (anonymous):

oh ok..in my book it says that (bsubn+1) has to be less than or equal to (bsubn) but also that (bsubn) has to equal 0 when the limit is taken...

OpenStudy (anonymous):

but \[\sum_{n=1}^{\infty}\frac{1}{n^3}\] also converges, so this series is not just convergent, it is absolutely convergent as well

OpenStudy (anonymous):

and i wasnt sure if the former was true

OpenStudy (anonymous):

yeah only need the terms go to zero

OpenStudy (anonymous):

ok, thanks satellite

OpenStudy (anonymous):

\[\sum\frac{(-1)^n}{n}\] for example converges, although not absolutely

OpenStudy (anonymous):

hell, even \[\sum\frac{(-1)^n}{\ln(n)}\] converges

OpenStudy (anonymous):

its converges conditionally right?

OpenStudy (anonymous):

the last two i wrote are only conditionally convergent, the one you posted is absolutely convergent, a stronger condition

OpenStudy (anonymous):

ok thanks you for your help! much appreciated

OpenStudy (anonymous):

yw